OpenStudy (anonymous):

Find the slope-intercept equation for the line with slope 5/7 and y-intercept (0,4) y = mx + b Find b: 4 = 5/7(0) + b 4 = 0 + b b = 4. Once we have m and b, substitute into y = mx+b. So what is our answer, Tracey?

OpenStudy (anonymous):

y = (5/7)x + 4 7y = 5x + 28

OpenStudy (anonymous):

Im confused now how would I put this on a chart

OpenStudy (anonymous):

Hi, Prewitt -- You can go from the given info DIRECTLY to y =5/7x + 4. If y-intercept is (0, 4) then b = 4! To graph it, put a dot at the point (0, 4) then from that point run over 7 units to the right and up 5 to the point (7, 9). Connect the dots.